# Path-specific effects for pulse-oximetry guided decisions in critical care

### v3

@@ Abstract
Identifying and measuring biases associated with sensitive attributes is a crucial consideration in healthcare to prevent treatment disparities. One prominent issue is inaccurate pulse oximeter readings, which tend to overestimate oxygen saturation for dark-skinned patients and misrepresent supplemental oxygen needs. Most existing research has revealed statistical disparities linking device measurement errors to patient outcomes in intensive care units (ICUs) without causal formalization. This study causally investigates how racial discrepancies in oximetry measurements affect invasive ventilation in ICU settings. We employ a causal inference-based approach using path-specific effects to isolate the impact of bias by race on clinical decision-making. To estimate these effects, we leverage a doubly robust estimator, propose its self-normalized variant for improved sample efficiency, and provide novel finite-sample guarantees. Our methodology is validated on semi-synthetic data and applied to two large real-world health datasets: MIMIC-IV and eICU. Contrary to prior work, our analysis reveals minimal impact of racial discrepancies on invasive ventilation rates. However, path-specific effects mediated by oxygen saturation disparity are more pronounced on ventilation duration, and the severity differs by dataset. Our work provides a novel pipeline for investigating potential disparities in clinical decision-making and, more importantly, highlights the necessity of causal methods to robustly assess fairness in healthcare.
@@ 1 Introduction
Bias in medical devices can perpetuate disparities by impacting healthcare decisions. For example, pulse oximeters tend to overestimate blood oxygen saturation for dark-skinned patients [45, 50, 43, 21, 42, 15]. This discrepancy can lead to ‘hidden hypoxemia,’ where a patient’s true arterial oxygen saturation ($\text{SaO}_{2}$) is dangerously low, despite a reassuringly higher peripheral $\text{SpO}_{2}$ reading from the oximeter. Figure 1 demonstrates this using eICU data [40], where red dots (‘H.H’) highlight instances where low $\text{SaO}_{2}$ (< 88%) is masked by higher $\text{SpO}_{2}$ (> 88%). Such hidden hypoxemia can result in underestimated supplemental oxygen needs and delayed clinical interventions [50, 14, 12].
@@
    We apply our framework to analyze race-based disparities mediated by oximeter discrepancies on invasive ventilation (rates and duration) in MIMIC-IV [22] and eICU [40].
@@ 2 Preliminaries
Notation. We use ($\mathbf{X}$, $X$, $\mathbf{x}$, $x$) to denote a random vector, variable, and their realized values, respectively. For a function $f(\mathbf{z})$, we use $\sum_{\mathbf{z}}f(\mathbf{z})$ to denote the summation/integration over discrete/continuous $\mathbf{Z}$. For a discrete vector $\mathbf{X}$, we denote $1[\mathbf{X}=\mathbf{x}]$ as an indicator function such that $1[\mathbf{X}=\mathbf{x}]=1$ if $\mathbf{X}=\mathbf{x}$ and 0 otherwise. $P(\mathbf{V})$ denotes a distribution over $\mathbf{V}$ and $P(\mathbf{v})$ a probability at $\mathbf{V}=\mathbf{v}$. We let $\mathbb{E}[f(\mathbf{V})]$ and $\mathbb{V}[f(\mathbf{V})]$ denote the mean and variance of $f(\mathbf{V})$ relative to $P(\mathbf{V})$, and $\|f\|_{P}\triangleq\sqrt{\mathbb{E}[(f(\mathbf{V}))^{2}]}$ as the $L_{2}$-norm of $f$ with $P$. We use $\widehat{f}-f=o_{P}(r_{n})$ if $\widehat{f}$ is a consistent estimator of $f$ having rate $r_{n}$, and $\widehat{f}-f=O_{P}(r_{n})$ if $\widehat{f}-f$ is bounded in probability at rate $r_{n}$. We will say $\hat{f}$ is $L_{2}$-consistent if $\|\hat{f}-f\|_{P}=o_{P}(1)$ and $\widehat{f}-f=O_{P}(1)$ if $\widehat{f}-f$ is bounded in probability. Let $\mathcal{D}\triangleq\{\mathbf{V}_{i}:i=1,\cdots,n\}$ denotes a set of $n$ samples. The empirical average of $f(\mathbf{V})$ over $\mathcal{D}$ is denoted as $\mathbb{E}_{\mathcal{D}}[f(\mathbf{V})]\triangleq(1/|\mathcal{D}|)\sum_{i:\mathbf{V}_{i}\in\mathcal{D}}f(\mathbf{V}_{i})$.
@@ 2.1 Related work
Much of the literature has focused on estimating NDE and NIE under single-mediator settings (e.g., [47, 11]). Motivated by presence of multiple mediators, contemporary research on path-specific analysis has broadened its scope to investigate effects transmitted along particular, often complex, paths within multi-mediator systems. For instance, effects acting through the colored path in Fig. 2(a) might be characterized by the nested counterfactual quantity $\mathbb{E}[Y_{x_{1},V_{x_{0},W_{x_{1}}}}]$ in statistics and algorithmic fairness [49, 30, 52, 6, 8]. Specifically, [30] developed a robust estimator for measuring path-specific effects transmitted through a particular mediator in the presence of other mediators.
@@ 3.1 Estimating path-specific effects
Under the SFMs in Fig.  2, the VDE $\mathbb{E}[Y_{x_{1}}]-\mathbb{E}[Y_{x_{1},V_{x_{0},W_{x_{1}}}}]$ is identifiable and is given as follows: $\mathbb{E}[Y_{x_{1}}]=\sum_{z}\mathbb{E}[Y\mid x_{1},z]P(z)$ and
@@
In this section, we focus on estimating the nested counterfactual term of the VDE, $\mathbb{E}[Y_{x_{1},V_{x_{0},W_{x_{1}}}}]$ represented in Eq. (2), since estimating $\mathbb{E}[Y_{x_{1}}]$ using the back-door adjustment [35] is well-known.
To construct the estimator, we will parameterize Eq. (2) using following:
@@
Given a sample $\mathcal{D}\overset{iid}{\sim}P$, the doubly robust VDE estimator $\hat{\psi}$ is constructed as follows:
@@
Following the analysis from [Theorem 4 in 25], we provide novel finite sample learning guarantees:
Theorem 2 (Finite Sample Analysis).
@@ 4.1 Synthetic data
Figure 2: Convergence of the relative error of causal queries using the canonical (left column) and self-normalized (right column) doubly robust estimator on a continuous synthetic setting (top row) and semi-synthetic eICU data (bottom row). The plots show the mean and 95% confidence interval over 100 bootstraps iterations. The estimates using self-normalization exhibit significantly lower variance.
@@ 4.3 Semi-synthetic eICU experiments
Following the setup in the synthetic experiment, we compute the relative error for $\mathbb{E}[Y_{x_{0}}]$, $\mathbb{E}[Y_{x_{1}}]$, $\mathbb{E}[Y_{x_{1},M_{x_{0}}}]$, and $\mathbb{E}[Y_{x_{1},V_{x_{0},W_{x_{1}}}}]$ across varying finite sample sizes in Figures 2(c) and 2(d). Similar to the synthetic experiments, our proposed self-normalized variant exhibits significantly lower variance in a healthcare-grounded setting with complex relationships between variables. This behavior is consistent with results using MIMIC-based semi-synthetic data, which we include in Appendix G. Additionally, we compare our estimators with state-of-the-art effect estimation baselines in Appendix I and demonstrate their robustness to imbalance between $x_{0}$ and $x_{1}$ subpopulations in Appendix J.
@@ 5 Discussion
In this work, we introduce the $V$-specific Direct Effect (VDE) to quantify racial biases in critical care decision-making that are mediated by measurement discrepancies from pulse oximeter devices. While previous studies have documented disparate outcomes for Black patients, our approach is the first to apply a path-specific causal framework to examine heterogeneity by race in a clinically actionable healthcare process. We develop a doubly robust estimator and a self-normalized variant for the VDE, provide favorable finite sample guarantees, and demonstrate strong empirical performance on both synthetic and semi-synthetic health data.
@@
For assessing clinical decision-making, our findings indicate that bias arising from pulse oximetry measurements primarily affects the duration of invasive ventilation rather than initiation of treatment. This pattern suggests that clinicians may be effectively integrating additional patient information beyond oxygen saturation when deciding whether to ventilate. Alternatively, the bias may manifest through pathways not captured within our current framework, such as through unmeasured variables or outcomes. Understanding the mechanisms underlying such bias remains an active area of research.
@@ A.4 Proof of Lemma 2
For any functional $(a_{0}(\mathbf{W}),b_{0}(\mathbf{W}))$ and $(a(\mathbf{W}),b(\mathbf{W})$, for any $\mathbf{W}\subseteq\mathbf{V}$, the following holds:
@@ Appendix C Additional data analysis
Figure 7: Distribution of patient baseline characters, (a) age and (b) sex, in the eICU and MIMIC-IV datasets, stratified by race.

### v4

@@ Abstract
Identifying and measuring biases associated with sensitive attributes is a crucial consideration in healthcare to prevent treatment disparities. One prominent issue is inaccurate pulse oximeter readings, which tend to overestimate oxygen saturation for dark-skinned patients and misrepresent supplemental oxygen needs. Most existing research has revealed statistical disparities linking device measurement errors to patient outcomes in intensive care units (ICUs) without causal formalization. This study causally investigates how racial discrepancies in oximetry measurements affect invasive ventilation in ICU settings. We employ a causal inference-based approach using path-specific effects to isolate the impact of bias by race on clinical decision-making. To estimate these effects, we leverage a doubly robust estimator, propose its self-normalized variant for improved sample efficiency, and provide novel finite-sample guarantees. Our methodology is validated on semi-synthetic data and applied to two large real-world health datasets: MIMIC-IV and eICU. Contrary to prior work, our analysis reveals minimal impact of racial discrepancies on invasive ventilation rates. However, path-specific effects mediated by oxygen saturation disparity are more pronounced on ventilation duration, and the severity differs across datasets. Our work provides a novel pipeline for investigating potential disparities in clinical decision-making and, more importantly, highlights the necessity of causal methods to robustly assess fairness in healthcare.
@@ 1 Introduction
Bias in medical devices can perpetuate disparities by impacting healthcare decisions. For example, pulse oximeters tend to overestimate blood oxygen saturation for dark-skinned patients [45, 50, 43, 21, 42, 15]. Such discrepancies can lead to ‘hidden hypoxemia,’ where a patient’s true arterial oxygen saturation ($\text{SaO}_{2}$) is dangerously low, despite a reassuringly higher peripheral $\text{SpO}_{2}$ reading from the oximeter. Figure 1 demonstrates this using eICU data [40], where red dots (‘H.H’) highlight instances where low $\text{SaO}_{2}$ (< 88%) is masked by higher $\text{SpO}_{2}$ (> 88%). Such hidden hypoxemia can result in underestimated supplemental oxygen needs and delayed clinical interventions [50, 14, 12].
@@
    We apply our framework to analyze race-based disparities mediated by oximeter discrepancies on invasive ventilation (rate and duration) in MIMIC-IV [22] and eICU [40].
@@ 2 Preliminaries
Notation. We use ($\mathbf{X}$, $X$, $\mathbf{x}$, $x$) to denote a random vector, variable, and their realized values, respectively. For a function $f(\mathbf{z})$, we use $\sum_{\mathbf{z}}f(\mathbf{z})$ to denote the summation/integration over discrete/continuous $\mathbf{Z}$. For a discrete vector $\mathbf{X}$, we denote $1[\mathbf{X}=\mathbf{x}]$ as an indicator function such that $1[\mathbf{X}=\mathbf{x}]=1$ if $\mathbf{X}=\mathbf{x}$ and 0 otherwise. $P(\mathbf{V})$ denotes a distribution over $\mathbf{V}$ and $P(\mathbf{v})$ a probability at $\mathbf{V}=\mathbf{v}$. We let $\mathbb{E}[f(\mathbf{V})]$ and $\mathbb{V}[f(\mathbf{V})]$ denote the mean and variance of $f(\mathbf{V})$ relative to $P(\mathbf{V})$, and $\|f\|_{P}\triangleq\sqrt{\mathbb{E}[(f(\mathbf{V}))^{2}]}$ as the $L_{2}$-norm of $f$ with $P$. We use $\widehat{f}-f=o_{P}(r_{n})$ if $\widehat{f}$ is a consistent estimator of $f$ having rate $r_{n}$, and $\widehat{f}-f=O_{P}(r_{n})$ if $\widehat{f}-f$ is bounded in probability at rate $r_{n}$. We will say $\hat{f}$ is $L_{2}$-consistent if $\|\hat{f}-f\|_{P}=o_{P}(1)$ and $\widehat{f}-f=O_{P}(1)$ if $\widehat{f}-f$ is bounded in probability. Let $\mathcal{D}\triangleq\{\mathbf{V}_{i}:i=1,\cdots,n\}$ denote a set of $n$ samples. The empirical average of $f(\mathbf{V})$ over $\mathcal{D}$ is denoted as $\mathbb{E}_{\mathcal{D}}[f(\mathbf{V})]\triangleq(1/|\mathcal{D}|)\sum_{i:\mathbf{V}_{i}\in\mathcal{D}}f(\mathbf{V}_{i})$.
@@ 2.1 Related work
Much of the literature has focused on estimating NDE and NIE under single-mediator settings (e.g., [47, 11]). Motivated by the presence of multiple mediators, contemporary research on path-specific analysis has broadened its scope to investigate effects transmitted along particular, often complex, paths within multi-mediator systems. For instance, effects acting through the colored path in Fig. 2(a) might be characterized by the nested counterfactual quantity $\mathbb{E}[Y_{x_{1},V_{x_{0},W_{x_{1}}}}]$ in statistics and algorithmic fairness [49, 30, 52, 6, 8]. Specifically, [30] developed a robust estimator for measuring path-specific effects transmitted through a particular mediator in the presence of other mediators.
@@ 3.1 Estimating path-specific effects
Under the SFMs in Fig. 2, the VDE $\mathbb{E}[Y_{x_{1}}]-\mathbb{E}[Y_{x_{1},V_{x_{0},W_{x_{1}}}}]$ is identifiable and is given as follows: $\mathbb{E}[Y_{x_{1}}]=\sum_{z}\mathbb{E}[Y\mid x_{1},z]P(z)$ and
@@
We focus on estimating the nested counterfactual term of the VDE, $\mathbb{E}[Y_{x_{1},V_{x_{0},W_{x_{1}}}}]$ represented in Eq. (2), since estimating $\mathbb{E}[Y_{x_{1}}]$ using the back-door adjustment [35] is well-known.
To construct the estimator, we will parameterize Eq. (2) using the following:
@@
Given a sample $\mathcal{D}\overset{i.i.d.}{\sim}P$, the doubly robust VDE estimator $\hat{\psi}$ is constructed as follows:
@@
Following the analysis from [Theorem 4 in 25], we provide novel finite-sample learning guarantees:
Theorem 2 (Finite-Sample Analysis).
@@ 4.1 Synthetic data
Figure 2: Convergence of the relative error of causal queries using the canonical (left column) and self-normalized (right column) doubly robust estimator on a continuous synthetic setting (top row) and semi-synthetic eICU data (bottom row). The plots show the mean and 95% confidence interval over 100 bootstrap iterations. The estimates using self-normalization exhibit significantly lower variance.
@@ 4.3 Semi-synthetic eICU experiments
Following the setup in the synthetic experiment, we compute the relative error for $\mathbb{E}[Y_{x_{0}}]$, $\mathbb{E}[Y_{x_{1}}]$, $\mathbb{E}[Y_{x_{1},M_{x_{0}}}]$, and $\mathbb{E}[Y_{x_{1},V_{x_{0},W_{x_{1}}}}]$ across varying sample sizes in Figures 2(c) and 2(d). Similar to the synthetic experiments, our proposed self-normalized variant exhibits significantly lower variance in a healthcare-grounded setting with complex relationships between variables. This behavior is consistent with results using MIMIC-based semi-synthetic data, which we include in Appendix G. Additionally, we compare our estimators with state-of-the-art effect estimation baselines in Appendix I and demonstrate their robustness to imbalance between $x_{0}$ and $x_{1}$ subpopulations in Appendix J.
@@ 5 Discussion
In this work, we introduce the $V$-specific Direct Effect (VDE) to quantify racial biases in critical care decision-making that are mediated by measurement discrepancies from pulse oximeter devices. While previous studies have documented disparate outcomes for Black patients, our approach is the first to apply a path-specific causal framework to examine heterogeneity by race in a clinically actionable healthcare process. We develop a doubly robust estimator and a self-normalized variant for the VDE, provide favorable finite-sample guarantees, and demonstrate strong empirical performance on both synthetic and semi-synthetic health data.
@@
For assessing clinical decision-making, our findings indicate that bias arising from pulse oximetry measurements primarily affects the duration of invasive ventilation rather than initiation of treatment. This pattern suggests that clinicians may effectively integrate additional patient information beyond oxygen saturation when deciding whether to ventilate. Alternatively, the bias may manifest through pathways not captured within our current framework, such as through unmeasured variables or outcomes. Understanding the mechanisms underlying such bias remains an active area of research.
@@ A.4 Proof of Lemma 2
For any functional $(a_{0}(\mathbf{W}),b_{0}(\mathbf{W}))$ and $(a(\mathbf{W}),b(\mathbf{W}))$, for any $\mathbf{W}\subseteq\mathbf{V}$, the following holds:
@@ Appendix C Additional data analysis
Figure 7: Distribution of patient baseline characteristics, (a) age and (b) sex, in the eICU and MIMIC-IV datasets, stratified by race.